### v3

# LLM-Guided Scenario-based GUI Testing

Abstract.
The assurance of mobile application (app) GUI has become increasingly important, as the GUI serves as the primary medium of interaction between users and apps. Although numerous automated GUI testing approaches have been developed with diverse strategies, a substantial gap remains between these approaches and the underlying app business logic. Most existing approaches focus on general exploration rather than the completion of specific testing scenarios, often resulting in missed coverage of critical functionalities. Inspired by the manual testing process, which treats business logic, driven testing scenarios as the fundamental unit of testing, this paper introduces an approach that leverages large language models (LLMs) to comprehend the semantics expressed in app GUIs and their contextual relevance to given testing scenarios. Building upon this capability, we propose ScenGen, a novel LLM-guided scenario-based GUI testing framework that employs a multi-agent collaboration mechanism to simulate and automate the phases of manual testing.
Specifically, ScenGen integrates five agents: the Observer, Decider, Executor, Supervisor, and Recorder. The Observer perceives the app GUI state by extracting and structuring GUI widgets and layouts, thereby interpreting the semantic information presented in the GUI. This information is then passed to the Decider, which makes scenario-driven decisions with the guidance of LLMs to identify target widgets and determine appropriate actions toward fulfilling specific testing goals. The Executor executes the decided operations on the app, while the Supervisor verifies whether the execution results align with the intended scenario completion, ensuring traceability and consistency in test generation and execution. Finally, the Recorder records the corresponding GUI operations into the context memory as a knowledge base for subsequent decision-making and concurrently monitors runtime bug occurrences. Comprehensive evaluations demonstrate that ScenGen effectively generates scenario-based GUI tests guided by LLM collaboration, achieving higher relevance to app business logic and improving the completeness of automated GUI testing.
Mobile App Testing, GUI Testing, Large Language Model, Image Understanding, Multi-Agent Framework
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXjournal: TOSEMjournalvolume: 1journalnumber: 1article: 1publicationmonth: 1ccs: Software and its engineering Software testing and debugging
1. Introduction
With the continuous advancement of mobile devices and mobile Internet technologies, mobile applications (apps) have experienced rapid and large-scale development (Yu et al., 2024b). The accelerated iteration cycles of modern app ecosystems continuously drive their functional and interface evolution. This phenomenon not only enriches user experiences and diversifies app types but also imposes new challenges and higher demands on mobile app quality assurance (Yu et al., 2024a). Among various quality assurance techniques, GUI testing has become indispensable, as the GUI represents the primary interaction layer between users and apps (White et al., 2019).
Manual GUI testing, which relies on human testers to explore and interact with the app GUI, remains an intuitive approach for identifying usability issues, interaction anomalies, and visual inconsistencies. However, this process is inherently labor-intensive and time-consuming, making it impractical for frequent app updates and continuous delivery scenarios (Su et al., 2017). Consequently, researchers and practitioners have increasingly turned to automated GUI testing to improve efficiency, scalability, and repeatability. A common paradigm in automated GUI testing is test script record and replay (Talebipour et al., 2021; Behrang and Orso, 2019), where test scripts are recorded from manual interactions and later executed automatically using frameworks such as Appium (Appium, [n. d.]). While this approach enhances execution efficiency, it still heavily depends on manual scripting and frequent maintenance. The tight coupling between scripts and specific GUI structures makes these tests brittle when the app evolves (Yoon et al., 2022), leading to high maintenance costs and poor adaptability to new versions or requirements.
To overcome these limitations, automated test generation techniques have emerged. These approaches automatically explore the app under test to generate test cases according to predefined exploration strategies, thereby reducing manual intervention and enhancing testing efficiency. Common strategies include random-based, model-based, and learning-based approaches. Random-based strategies (mon, 2024; Machiry et al., 2013) rely on pseudo-random event generation to explore diverse GUI states, effectively uncovering crashes and inconsistencies. However, they often lack direction and scenario awareness. Model-based strategies (Su et al., 2017; Mao et al., 2016) improve exploration precision by constructing GUI models through static or dynamic analysis, which are then used to guide event generation. Yet, such models may struggle to capture semantic intent or user-level goals. Learning-based strategies (Yu et al., 2024b; Pan et al., 2020; Adamo et al., 2018) leverage advanced AI techniques, such as reinforcement learning, to navigate large and complex GUI state spaces. Although these approaches achieve higher adaptability, they still primarily optimize quantitative coverage metrics rather than aligning with app business logic.
Most existing automated GUI testing approaches are designed to maximize code coverage (Yu et al., 2024a), which, while useful, introduces several limitations. First, generating continuous, logically coherent event sequences that correspond to realistic user behaviors remains difficult (Yu et al., 2024b). As a result, these approaches fail to reproduce complex task flows that represent actual usage patterns. Second, automatically generating contextually appropriate inputs to trigger targeted app responses is still a major challenge (Liu et al., 2024b), leading to superficial interaction exploration that neglects semantic completeness. Third, using code coverage as the primary evaluation metric inadequately reflects the true quality of testing, since it measures execution breadth but not the functional or experiential depth of user interactions (Yu et al., 2024a). The root cause of these limitations lies in the lack of scenario-level understanding of app functionalities (Yu et al., 2024a, 2023). Existing automated approaches focus on state transitions and event diversity rather than on testing scenarios that mirror real user intentions and business workflows. Without such scenario comprehension, testing processes are prone to overlooking essential functionalities, resulting in incomplete validation and missed defects (Yu et al., 2025).
Therefore, to advance automated GUI testing, there is a pressing need for scenario-aware approaches that can bridge the gap between low-level GUI exploration and high-level business logic. Such approaches should be capable of understanding the semantics embedded in GUI widgets and layouts, reasoning about user intentions, and generating coherent test sequences that align with realistic usage scenarios. In essence, automated GUI testing needs to evolve from event-level exploration toward LLM-guided, human-like reasoning, where each test scenario is treated as a goal-oriented task consistent with the way human testers think and operate. This motivates the design of our proposed approach, ScenGen, which integrates LLM–based reasoning into scenario-based GUI testing to achieve semantically meaningful, context-aware, and traceable test generation.
In response to the aforementioned challenges in automated GUI testing, this paper introduces ScenGen, a novel scenario-based automated GUI testing framework driven by large language models (LLMs). ScenGen is designed to bridge the gap between GUI-level exploration and scenario-level reasoning by enabling LLMs to understand both the visual semantics of app GUIs and the contextual intent of testing scenarios. To achieve this, ScenGen leverages multi-modal LLMs for semantic perception and a multi-agent collaborative architecture for intelligent test generation.
The design of ScenGen is inspired by the classic OODA Loop decision-making model: Observe, Orient, Decide, and Act, originally proposed for dynamic decision processes in complex environments (Boyd, 1987). This model aligns well with the iterative nature of GUI testing, where perception, reasoning, execution, and feedback occur cyclically. Drawing from this intuition, ScenGen conceptualizes automated GUI testing as an iterative human-like process consisting of five stages: observation, decision-making, execution, result inspection and feedback, and recording. Each stage corresponds to a specific cognitive or operational phase in manual testing, and together they form a closed decision loop that continuously refines the testing process. To operationalize this design, ScenGen employs a set of specialized agents, Observer, Decider, Executor, Supervisor, and Recorder, supported by a shared context memory that maintains the overall testing context. Each agent encapsulates distinct reasoning and functional capabilities, enabling modularity, interpretability, and traceability throughout the testing process.
The Observer acts as the perceptual component of ScenGen. It identifies and analyzes GUI widgets and layouts, addressing the inherent diversity and complexity of mobile GUIs. Accurate GUI recognition is fundamental for formulating effective testing actions. By combining computer vision and multi-modal LLM techniques, the Observer not only detects GUI components and their attributes but also interprets their semantic roles within the interface (e.g., distinguishing between navigation, input, or functional widgets). This design ensures that subsequent decision-making is grounded in a comprehensive and semantically rich understanding of the current app state.
Building upon the output of Observer, the Decider functions as the reasoning and planning agent. It interprets the recognized GUI elements in the context of the target testing scenario and generates a detailed action plan for test execution. The design of Decider enables independent decision-making by leveraging the scenario’s semantic goal, rather than random or purely structural exploration, to determine which widgets to interact with and how. This approach provides a logical foundation for goal-oriented testing, ensuring that generated operations align with business logic and the intended scenario objectives. Moreover, by decoupling decision-making from perception and execution, Decider enhances both the flexibility and generality of ScenGen, allowing it to adapt to diverse apps and scenarios.
The Executor is responsible for executing the planned operations on the app. By strictly following the action plan generated by Decider, the Executor ensures consistency, traceability, and repeatability in testing. Separating execution from decision-making also prevents cascading errors caused by incomplete reasoning or misinterpretation during planning. This modular separation mirrors human testing behavior, where decisions are reasoned out before actual interaction, thereby enhancing testing reliability.
The Supervisor serves as the feedback and validation agent. It inspects whether the executed operations conform to the intended plan and verifies whether the intermediate or final app states satisfy the requirements of the target scenario. These requirements are derived from two sources: explicit scenario descriptions provided at the start of exploration and implicit semantic understanding embedded within the LLMs. By formalizing this verification process, Supervisor allows ScenGen to identify deviations, provide corrective feedback, and maintain a consistent alignment between generated tests and scenario objectives. This continuous inspection and correction mechanism prevents error propagation and ensures that the generated tests remain semantically faithful to the intended user flows.
The Recorder is responsible for recording the complete interaction history, including GUI states, executed actions, and intermediate feedback, into the context memory, which serves as a persistent context memory shared among agents. This design choice enables iterative learning and context-aware decision-making across multiple testing cycles. The recorded context allows ScenGen to trace the reasoning behind past actions, reuse relevant knowledge, and dynamically adjust future exploration strategies. Additionally, Recorder monitors runtime logs and execution traces to identify potential failures or crashes, ensuring that discovered bugs are systematically documented. Once the recording is complete, the next iteration begins until the target scenario has been fully explored and validated.
Through this multi-agent, LLM-guided collaboration, ScenGen simulates the human tester’s reasoning process, observing, interpreting, deciding, acting, and learning iteratively, to achieve semantically coherent, scenario-oriented GUI testing.
The main contributions of this paper can be summarized as follows:
  • We propose ScenGen, a novel LLM-guided, scenario-based GUI testing framework that integrates multi-agent collaboration with GUI semantic understanding. The framework enables automated test generation aligned with high-level testing scenarios rather than low-level event exploration.
  • We design a GUI understanding and widget localization mechanism that combines computer vision and multi-modal LLMs to achieve comprehensive GUI comprehension and support semantic test generation.
  • We construct an app benchmark dataset containing ten representative testing scenarios and empirically evaluate ScenGen. The results demonstrate its effectiveness and efficiency in generating semantically meaningful, scenario-based GUI tests.
The replication package is available at https://github.com/iGUITest/ScenGen, and additional implementation details of ScenGen are provided at https://sites.google.com/view/scengen.
2. Background & Motivation
In this section, we first introduce the core concept underlying our work, the testing scenario, which serves as the foundation of the proposed approach. Then, we discuss the background and motivation for employing LLMs in the context of GUI testing.
2.1. Testing Scenario
A testing scenario represents a specific functionality or task within the app that encapsulates a complete and self-consistent business logic from the perspective of end users. Conceptually, a testing scenario reflects how users interact with the app through a coherent sequence of GUI operations to achieve a concrete goal, such as logging in, completing a purchase, or submitting a form. Each scenario corresponds to a meaningful user workflow and can be viewed as a semantic unit of testing that connects interface-level actions with business-level objectives.
In real-world app usage, end users interact with mobile apps through a series of GUI operations, such as tapping, swiping, or text input, to accomplish a specific goal. These operation sequences form logical patterns that can be abstracted into corresponding testing scenarios. For example, the “user login” scenario may include operations like entering credentials, tapping the login button, and verifying access to the home page. Understanding such scenarios is critical for testing app functionality from a user-centric perspective, as it enables testers to validate whether the app behaves as intended under realistic usage contexts. In the practice of manual GUI testing, developers and testers typically organize test cases around distinct testing scenarios. This organization allows them to design test cases more systematically, ensuring that all major functionalities and user flows are adequately exercised. Testing scenarios thus serve as a bridge between app functionality and user behavior, facilitating an intuitive and comprehensive verification of the app behavior.
However, despite its practical value, the concept of testing scenarios is largely neglected in existing automated GUI testing approaches. Current methods often emphasize event-level exploration, code coverage, or state-space traversal without explicitly modeling the app’s business logic or user goals (Yu et al., 2024a). As a result, these approaches frequently generate fragmented or redundant test sequences that lack semantic coherence. To bridge this gap, there is a need for testing methodologies that can explicitly model and reason about testing scenarios, recognizing not only how to interact with GUI elements but also why those interactions matter in achieving specific business objectives. This observation provides the conceptual motivation for our proposed framework, ScenGen, which introduces LLM-guided, scenario-based GUI testing to integrate high-level scenario understanding with automated test generation.
2.2. Motivating Examples
In this section, we present several real-world examples that illustrate the challenges faced by existing automated GUI testing techniques and motivate the design of ScenGen. These examples demonstrate why incorporating the semantic reasoning capabilities of LLMs and grounding test generation in testing scenarios is essential for improving testing completeness and relevance.
Figure 1. Motivating Examples
2.2.1. Hidden Function Triggering
Modern app GUIs often employ minimalist design principles, grouping related functionalities behind a single entry to maintain interface simplicity. These hidden widgets are typically accessed through expandable menus or contextual icons. As shown in Fig. 1 (a), in a note-taking app, the widget for creating a “text note” is nested within a hidden menu that appears only after clicking the plus icon in the bottom-right corner. If the user clicks outside the menu, the panel automatically closes. Such designs pose a significant challenge for automated GUI testing approaches that rely on surface-level widget exploration or static state transitions, as they may fail to discover the hidden functional paths. During the exploration of the “create a new note” scenario, LLMs can reason about the GUI’s semantic layout and infer that the plus icon is likely associated with content creation. This enables the Observer and Decider agents in ScenGen to proactively explore the hidden menu and trigger the corresponding functionality, effectively mimicking human intuition in navigating dynamic and context-dependent interfaces.
2.2.2. Domain Knowledge Dependency
In many cases, effective GUI testing depends not only on interface exploration but also on domain-specific knowledge relevant to the app’s functionality. As illustrated in Fig. 1 (b), although calculator apps have simple GUIs, the correctness of their functionality depends on mathematical boundary conditions and logical rules (e.g., division by zero, large number handling). Conventional automated testing methods struggle to identify and generate meaningful test inputs that capture such domain constraints, often limiting their ability to validate correctness comprehensively. By contrast, LLMs possess the capacity to combine GUI understanding with domain reasoning. Within ScenGen, the LLM-driven agents can interpret the purpose of calculator widgets, associate them with mathematical operations, and generate scenario-based test sequences that cover critical boundary and functional conditions. This integration of domain knowledge allows the testing process to move beyond surface-level interactions and focus on semantic correctness, achieving deeper functional coverage.
2.2.3. Complicated GUI Event Generation
Many app functionalities are realized through multi-step interaction sequences that reflect the underlying business logic. For instance, as depicted in Fig. 1 (c), a typical login process involves sequential operations such as entering credentials, selecting options, and submitting the form. Each step is temporally dependent and semantically constrained. For example, the “submit” button should only be triggered after valid input fields are filled. Traditional automated GUI testing approaches that rely on random or model-based exploration often fail to preserve such logical dependencies, producing invalid or incomplete event sequences. In contrast, ScenGen leverages LLMs to reason about the intended workflow and maintain contextual continuity across operations. By understanding the purpose of each widget and its relation to the overall scenario goal, LLM agents can generate coherent and meaningful event sequences, such as inputting valid user data followed by appropriate form submission, thus aligning the testing process with realistic user behavior.
3. Methodology
ScenGen comprises multiple collaborative components, as illustrated in Fig. 2, including the context memory, Observer, Decider, Executor, Supervisor, and Recorder. These components jointly form an iterative process that integrates perception, reasoning, execution, verification, and feedback to achieve scenario-based GUI testing. Among them, several components, such as the Observer, Decider, and Supervisor, are powered by the LLMs to provide high-level semantic understanding and decision reasoning to ensure reliable and traceable test execution and validation.
Specifically, the Observer perceives and interprets the app GUI by detecting widgets and extracting their semantic attributes, enabling the system to understand the current interface context. Building upon this perception, the Decider performs reasoning and planning, leveraging LLM capabilities to align GUI interactions with the target testing scenario and business logic. The Executor then executes the planned operations on the app, ensuring repeatability and traceability while decoupling action execution from decision-making to reduce the risk of cascading failures. Next, the Supervisor verifies whether the executed operations achieve the intended scenario objectives and provides feedback for correction when discrepancies occur. Finally, the Recorder records all operations, GUI states, and inspection outcomes into the shared context memory, preserving contextual data to support consistent and informed decision-making in subsequent iterations.
Through the collaboration of these LLM-driven and rule-based components, ScenGen progressively refines its understanding of app functionality and testing intents until all target scenarios are comprehensively explored and validated. To ensure reproducibility and transparency, all prompt templates used for agent interactions are publicly available in our online resources.
Figure 2. ScenGen Framework
3.1. Context Memory
The context memory is designed to maintain continuity and coherence throughout the iterative testing process, ensuring that ScenGen can generate logically consistent scenario-based GUI tests. However, large language models (LLMs) are inherently stateless, lacking the ability to retain contextual information between interactions. This limitation poses a major challenge for multi-step GUI testing, where each decision depends on accumulated knowledge of previous actions, app states, and scenario progress. Without persistent context, the testing process risks producing inconsistent or redundant behaviors. To overcome this issue, ScenGen introduces a structured context memory inspired by human memory theory (Baddeley, 1992). The context memory acts as a central repository for storing and managing diverse types of information generated during testing, thereby supporting reasoning, decision-making, and feedback across agents. Specifically, it is divided into three components: long-term memory, which stores static information such as device configuration, app metadata, and testing scenarios; working memory, which maintains dynamic information including current testing goals, executed actions, and agent interactions; and short-term memory, which records transient GUI states and recognition results. This hierarchical organization enables seamless information flow across iterations, ensuring that testing decisions remain contextually grounded and progressively adaptive.
Long-term Memory. This memory stores static and global information that remains consistent throughout the entire testing campaign. It includes details about the testing environment (e.g., device information, operating system version), app metadata (e.g., app name, package identifier, version), and descriptions of the testing scenarios. These data are organized in a structured format and provide stable background knowledge for the LLM-based agents. The long-term memory ensures that agents can interpret GUI semantics and scenario goals in the correct app context, eliminating the need for redundant reinitialization across iterations or app sessions.
Working Memory. The working memory retains dynamic but persistent information relevant to the current testing session. It contains the target testing scenario, the current app under test, the history of executed operations, intermediate outcomes, and the recent interaction history with the LLMs. Unlike long-term memory, working memory evolves continuously as the test progresses, reflecting changes in GUI states and user-like actions. Each agent accesses this memory to maintain consistency and traceability. For example, the Decider consults it to decide the next operation based on past actions, while the Supervisor refers to it when verifying whether a sequence of actions aligns with the intended scenario. The structured storage enables the system to efficiently retrieve relevant records for reasoning and prompt generation.
Short-term Memory. The short-term memory captures transient and immediate data associated with the most recent interactions. It stores the current GUI state, the preceding GUI state, and the results of widget recognition performed by the Observer. Such information is frequently updated as the app interface evolves during testing. The app states are represented primarily as GUI screenshots, augmented with detected widget attributes and layout structures. The short-term memory enables quick contextual reasoning, for example, allowing the Decider to compare the current GUI against the previous one to determine whether the intended operation succeeded or whether a new state transition occurred.
To improve efficiency and prevent redundant exploration, ScenGen employs a GUI state comparison mechanism within the short-term memory. Two GUI states are considered identical if their visual and structural representations differ only in trivial details (e.g., minor animations, timestamps, or non-functional layout shifts). This comparison leverages perceptual hashing and structural similarity metrics to identify equivalent states. When an identical state is detected, the system skips re-exploration and updates only the relevant context, significantly reducing unnecessary operations and improving testing efficiency.
Overall, the context memory forms the cognitive foundation of ScenGen iterative process. It is dynamically updated during widget recognition, action execution, and test recording, while decision-making and validation stages rely on its stored content to formulate prompts, reason about app behavior, and ensure consistency across testing iterations. This design enables the agents in ScenGen to emulate human-like continuity and memory-driven reasoning during automated scenario-based GUI testing.
3.2. Observer
The Observer is responsible for perceiving the app’s GUI and constructing a structured representation of its current state to support subsequent decision-making and testing operations. However, accurately understanding GUI states presents a dual challenge: traditional computer vision (CV) techniques offer precise pixel-level detection but lack semantic comprehension, while large language models (LLMs) possess strong reasoning abilities yet struggle with fine-grained localization. To overcome this limitation, ScenGen adopts a hybrid perception design that combines traditional CV algorithms with multi-modal LLM reasoning. The CV component extracts and localizes widgets through visual analysis, while the LLM interprets their semantic roles and contextual relationships within the interface. This integration enables the Observer to achieve both spatial precision and semantic depth, producing a comprehensive, human-like understanding of the GUI state essential for accurate, scenario-driven test generation.
The widget recognition process primarily targets GUI screenshots and proceeds in three main stages. First, CV algorithms are employed to identify potential GUI widgets within the GUI image. Second, optical character recognition (OCR) algorithms are applied to extract textual elements from the same screenshot, capturing words, phrases, and textual labels. These two stages provide the foundational visual and textual components for the subsequent fusion stage. Finally, the recognition results are integrated and refined to generate the final widget set, ensuring that both visual completeness and semantic coherence are preserved.
During this fusion and filtering stage, several refinements are applied. We merge adjacent text fragments based on spatial proximity to form coherent textual widgets and exclude unreasonable GUI detections, such as duplicated text regions, status-bar icons (e.g., battery indicators or signal symbols), or objects that are excessively large (encompassing multiple widgets) or too small (noise artifacts). Additionally, related graphical and textual widgets are merged into unified widget entities when they jointly represent the same user-interaction target (e.g., an icon with a label underneath). The merging is guided by a predefined distance threshold, following prior empirical settings (Yu et al., 2024b, a).
It is important to note that ScenGen adopts a vision-based recognition strategy rather than relying solely on extracting widget information from GUI layout files. This decision is motivated by practical limitations: widgets embedded in Canvas objects or hybrid H5 pages may be inaccessible through layout parsing (Yu et al., 2025, 2024a, 2024b). Directly identifying widgets from rendered screenshots ensures a “what-you-see-is-what-you-get” representation of the GUI, aligning the perceived interface with the user’s actual view and preserving fidelity in testing.
After obtaining the final widget recognition results, the Observer invokes a multi-modal LLM instance to perform semantic interpretation of the GUI state. The LLM analyzes visual layouts and textual cues to infer the roles and relationships of widgets (e.g., recognizing navigation buttons, input fields, or functional icons). This semantic understanding enriches the purely structural detection results from the CV stage, enabling subsequent agents (Decider, Supervisor) to reason about GUI intent and business logic during scenario-based testing.
Finally, all recognized widgets are visualized by overlaying bounding boxes and IDs on the original screenshot. This annotated image, along with the structured widget list and the raw GUI image, is stored in the short-term memory of ScenGen. This visual-structural representation not only facilitates cross-agent communication but also serves as a crucial input for LLM-based reasoning and action planning in later testing stages.
3.3. Decider
The Decider is responsible for determining the next operation during scenario-based GUI test generation, guiding the testing process toward fulfilling the target scenario. In ScenGen, test generation is modeled as an iterative decision-making process that begins from the app’s initial state and advances through a sequence of context-aware actions until the testing goal is achieved. However, this process presents two key challenges: first, accurately reasoning about the next action requires understanding both the current GUI state and the overall business logic of the target scenario; second, LLMs, while capable of semantic reasoning, may struggle to align their decisions with precise, executable GUI operations. To address these issues, ScenGen employs an LLM-guided decision mechanism that transforms GUI understanding and scenario semantics into concrete, stepwise operations. Each iteration evaluates the current app state, the history of executed actions, and the predefined scenario context to infer the most appropriate next operation through LLM reasoning. This design enables the Decider to bridge abstract scenario understanding and concrete GUI interactions, ensuring that every decision aligns logically with the scenario progression and contributes effectively to scenario completion. Every action decision is derived as an inference mapping generated by an LLM, formulated as follows:
(1) $(T,S,L_{op},P_{curr},W_{curr})\stackrel{{\scriptstyle LLM}}{{\longrightarrow}}op$
Structure of Target Scenarios. A target testing scenario $S$ represents a functional goal expressed through natural language, describing the intended user operation sequence or outcome. Each scenario is composed of (1) a concise textual description (e.g., “log into the app using valid credentials”) and (2) optional contextual information such as the expected starting page or relevant GUI screenshots. This natural language definition allows the LLM-based Decider to interpret the high-level intent and decompose it into a sequence of coherent GUI actions. The flexibility of scenario descriptions enables the framework to generalize across diverse apps while still grounding reasoning in the specific GUI context provided by the Observer.
Compared to traditional LLMs, multi-modal LLMs can directly process GUI screenshots, enabling visual-semantic reasoning and action planning. Thus, ScenGen employs raw GUI images and utilizes multi-modal LLMs to make visual decisions directly. However, this design introduces a cognition–execution gap: while LLMs excel at reasoning about what should be done, they lack the ability to locate precise pixel-level widget positions required for action execution (Liu et al., 2024b). To address this, ScenGen integrates multi-modal reasoning with traditional CV-based widget localization. The decision process is thus divided into two stages, logical decision-making and widget localization, bridging high-level reasoning and low-level action feasibility.
Figure 3. LLM Interaction of Decider
3.3.1. Logical Decision-making
The logical decision-making stage leverages the reasoning capabilities of multi-modal LLMs to analyze GUI images, interpret the current state, and decide the next logical operation. By combining the action type set $T$, the target scenario $S$, and the list of previously executed actions $L_{op}$, the LLM determines the next appropriate action $op^{\prime}$:
(2) $(T,S,L_{op},P_{curr})\stackrel{{\scriptstyle LLM}}{{\longrightarrow}}op^{\prime}$
The decision output $op^{\prime}$ contains: (1) the type of action (e.g., click, input, scroll), (2) a natural language description of the action intent, (3) a description of the target widget, (4) the input text (for input actions), and (5) the scroll direction (for scroll actions). For efficiency, ScenGen merges the implicit “focus click” operation required for input actions into a single composite action, avoiding redundant planning steps. Fig. 3 (a) illustrates this process.
Handling Complex Actions. Some scenarios require executing compound actions, i.e., coordinated sequences of interactions across multiple widgets (e.g., entering a username and password, then clicking a login button). The Decider manages these cases through dynamic action composition, in which the LLM decomposes the high-level intent into smaller, coherent sub-operations that maintain contextual continuity. Each sub-action is generated sequentially, guided by the scenario context and previous decisions stored in the context memory. This design enables flexible reasoning over multi-step workflows without requiring predefined action templates.
Text Value Generation. For input actions, the text content is dynamically produced according to the scenario context. The Decider first interprets semantic cues from the scenario description (e.g., “enter username”, select a realistic username). If the scenario provides explicit textual values, these are directly used; otherwise, the LLM generates contextually plausible text based on general knowledge or app-specific hints extracted from the GUI (e.g., placeholders, labels). This approach ensures that input data remain both semantically valid and consistent with user-intent reasoning, enabling realistic test execution.
3.3.2. Target Widget Location
After logical decision-making, the action intent and the target description are available, but precise pixel-level widget coordinates are still unknown. To enable actionable execution, the Decider performs widget localization, refining the decision into an executable operation:
(3) $(op^{\prime},W_{curr})\stackrel{{\scriptstyle LLM}}{{\longrightarrow}}op$
Widget Matching. Widget matching aligns the abstract target description from the decision stage with concrete GUI elements identified by the Observer. To ensure contextual coherence, ScenGen provides the LLM with visualized widget recognition results (annotated bounding boxes and IDs) as prompts, enabling it to correlate semantic intent with the correct widget on the interface. This direct comparison between annotated and raw screenshots ensures accurate widget identification, as shown in Fig. 3 (b).
Widget Prediction. In cases where certain widgets are visually indistinct or missing from recognition outputs (e.g., due to blurred boundaries or dynamic layouts), ScenGen employs a widget prediction mechanism. Using the relative positions and spatial relationships of nearby recognized widgets, the framework estimates the most probable location of the missing target, generating a “virtual widget” to maintain logical flow. This ensures robustness against incomplete GUI detection, as illustrated in Fig. 3 (c).
Widget Location Adjustment. Before executing input actions, ScenGen adjusts click positions to match functional regions. For example, if the target “To:” label is identified during email composition, the system shifts the click focus to the adjacent text box, which is the actual interactive element. This adjustment guarantees that interactions correspond to functional widgets rather than purely decorative or descriptive elements (Yu et al., 2024a). Fig. 3 (d) illustrates such adjustment.
3.3.3. Self-Correction
When a recent decision fails state transition verification, indicating that the intended effect was not achieved, the self-correction mechanism is triggered. This mechanism includes two steps: cause analysis and decision adjustment. The cause analysis identifies one of three common error types: (1) logical decisions were correct, but widget localization failed; (2) logical decisions were reasonable, but required preconditions (e.g., navigation steps) were missed; or (3) the logical decision itself deviated from scenario intent.
Based on the identified cause, the correction is applied in the corresponding dialogue context. For widget localization errors, the context of the localization prompt is reused to request revised reasoning. For logical or contextual errors, the dialogue from the decision-making phase is revisited, and the LLM is prompted to produce an adjusted plan. Through this iterative feedback loop, ScenGen ensures that each decision advances testing progress coherently and aligns with the target scenario. An example of this self-correction process is shown in Fig. 3 (e).
3.4. Executor & Recorder
The Executor and Recorder in ScenGen are designed to ensure the stability, traceability, and reproducibility of test execution. These modules are responsible for translating the high-level action plans generated by preceding agents into concrete GUI interactions and for recording the corresponding results and runtime information. This separation between reasoning (handled by LLM-based agents) and execution (handled by rule-based modules) prevents error propagation from uncertain LLM outputs and guarantees consistent interaction with the app under test.
ScenGen supports a wide range of common GUI interaction operations, including click, input, scroll, back, and others, which together cover the majority of user actions encountered during mobile app testing. The execution process begins with the Executor parsing the structured action decision produced by the Decider. It then invokes system-level interfaces (e.g., Android ADB commands) to perform the specified operation on the device. After execution, the Recorder logs the outcomes, including executed actions, GUI states, and system responses, into the working memory of the context memory, ensuring that subsequent iterations have full context for reasoning and verification. Additionally, the Recorder monitors runtime logs to detect anomalies such as app crashes or unhandled exceptions, which are recorded as potential defects during exploration.
Each operation type may require different parameters during the parsing phase. For click actions, the target widget must be specified. Input operations require both the target widget and the text content, as well as information about the input’s positional relationship to the widget. Scroll operations demand a defined scroll direction and, optionally, a target region (defaulting to the entire screen if unspecified). Notably, input operations are executed as a two-step process: first, performing a click to focus on the input field, followed by the text entry. The back operation, by contrast, requires no additional context parameters.
By implementing execution and recording as deterministic processes, ScenGen ensures that every GUI interaction is repeatable and verifiable, forming a reliable foundation for multi-agent collaboration. This design not only strengthens traceability across test iterations but also facilitates accurate post-analysis of failures and system behaviors, supporting robust and transparent scenario-based GUI testing.
3.5. Supervisor
The Supervisor is responsible for verifying the app state after each executed action, ensuring that the operation outcomes align with expectations before the next decision is made. Accurate verification is essential to maintain consistency between scenario intent and actual app behavior during iterative test generation. However, GUI state validation in mobile apps is inherently challenging due to dynamic page transitions, asynchronous loading behaviors, and diverse feedback mechanisms. Without robust verification, minor visual delays or hidden transitions may lead to misjudged states, erroneous decisions, or premature scenario termination. To address these challenges, the Supervisor employs a three-stage verification mechanism consisting of real-time loading verification, state transition verification, and test completion verification. This design allows the system to first ensure the app has reached a stable state, then confirm that state changes correspond to the expected outcomes, and finally determine whether the target testing scenario has been completed. By continuously validating and correcting deviations, the Supervisor provides a feedback loop that enhances system reliability, ensuring accurate and traceable progress throughout scenario-based test execution.
Figure 4. LLM Interaction of Supervisor
Real-time Loading Verification. The execution of a test operation may temporarily place the app into an unstable state, such as a loading or transition phase triggered by a network request. In such cases, premature decision-making may lead to misinterpretation of the app state or the early triggering of self-correction. While a fixed waiting time could mitigate this issue, it is difficult to generalize and often harms testing efficiency. Instead, ScenGen leverages the visual understanding capability of a multi-modal LLM to dynamically determine whether the app has completed loading. By analyzing the current GUI image, the Supervisor detects loading indicators or transitional states and adjusts the waiting duration adaptively. This adaptive waiting strategy improves both the reliability and efficiency of the testing process. An example of this is shown in Fig. 4 (a).
State Transition Verification. After ensuring that the app is stable, the next step is to verify whether the most recent operation has produced the intended state transition. The Supervisor uses a multi-modal LLM to compare the current GUI $P_{curr}$ with the previous GUI $P_{prev}$ under the current operation $op_{latest}$ and target scenario $S$, as expressed by
(4) $(S,op_{latest},P_{curr},P_{prev})\stackrel{{\scriptstyle LLM}}{{\longrightarrow}}result$
If the verification result indicates failure, an additional comparison between $P_{curr}$ and $P_{prev}$ is conducted to determine whether any actual change has occurred:
(5) $(P_{curr},P_{prev})\stackrel{{\scriptstyle LLM}}{{\longrightarrow}}result$
This step distinguishes between legitimate state transitions and irrelevant or automatic changes, such as clock updates, animated banners, or transient notifications. When no observable or meaningful effect is detected, ScenGen categorizes the operation as a non-effective action. In such cases, the system initiates a self-correction mechanism that re-evaluates the previous decision and generates an alternative operation targeting the same functional goal.
In situations where the executed operation produces unintended side effects, for example, navigating to an incorrect page or triggering an unwanted dialog, the Supervisor identifies the deviation through semantic comparison of the GUI state. The framework then performs a contextual rollback: it executes a predefined reverse action (e.g., a “back” operation) or reverts to the last valid state stored in the context memory. Fig. 4 (b) provides an example of this verification process. This design ensures that the exploration process remains consistent with the target scenario and prevents error accumulation in subsequent iterations. When the GUI effect is brief or difficult to capture (e.g., a short-lived popup), the Supervisor records a sequence of screenshots during the verification window and uses the LLM to detect temporal changes, ensuring that short-duration effects are properly recognized.
Test Completion Verification. The final stage determines whether the testing process has reached the expected endpoint for the current scenario. Because the endpoint cannot be predetermined, the Supervisor performs real-time evaluation during exploration to detect completion conditions. Leveraging a multi-modal LLM, it integrates information from the target scenario $S$, the list of executed operations $L_{op}$, and both the current and previous GUI screenshots ($P_{curr}$ and $P_{prev}$) to assess progress:
(6) $(S,L_{op},P_{curr},P_{prev})\stackrel{{\scriptstyle LLM}}{{\longrightarrow}}result$
If the LLM determines that the functional goal of $S$ has been achieved or the testing path sufficiently covers the intended scenario, the testing task is automatically terminated. Conversely, if an incomplete state is detected, the system continues the exploration process, guided by the Decider. This ensures that ScenGen terminates testing precisely when the target functionality has been comprehensively validated. An illustrative example is shown in Fig. 4 (c).
Through this multi-level verification and correction mechanism, the Supervisor guarantees that the app state transitions are accurately recognized, transient or spurious effects are properly handled, and unintended side effects are mitigated. This design ensures robust, context-aware supervision throughout the entire scenario-based testing process.
4. Empirical Evaluation
4.1. Experiment Settings
To comprehensively evaluate the performance and practicality of ScenGen, we design five research questions (RQs) that assess its effectiveness, efficiency, and real-world applicability:
  • RQ1: How effective is the scenario-based test generation of ScenGen?
  • RQ2: How effective is the bug detection capability of ScenGen?
  • RQ3: How effective is the widget localization of ScenGen?
  • RQ4: How effective is the logical decision-making of ScenGen?
  • RQ5: How efficient is the scenario-based test generation of ScenGen?
These research questions jointly provide a comprehensive assessment of ScenGen from multiple perspectives. Specifically, RQ1 investigates the overall quality of the generated tests, verifying the ability of ScenGen to produce meaningful and scenario-aligned testing sequences. RQ2 assesses the framework’s effectiveness in identifying real-world software bugs, demonstrating its practical utility. RQ3 measures the precision of widget localization conducted by the Observer, focusing on accurate visual grounding within GUI states. RQ4 evaluates the accuracy of logical decision-making by the Decider, assessing the LLM’s ability to produce semantically valid and contextually relevant actions. Finally, RQ5 examines the computational and temporal efficiency of test generation. Together, these RQs provide a multi-dimensional understanding of the capability and robustness of ScenGen.
LLM Configuration. We employ the GPT-4V model (et al., 2024) as the LLM instance in ScenGen, leveraging its advanced visual understanding capabilities for reasoning about GUI layouts and scenario semantics. The model is accessed through the gpt-4-vision-preview API, which enables direct visual processing of app screenshots. GPT-4V is chosen for its strong performance in multi-modal reasoning tasks, its ability to interpret complex visual contexts in combination with textual prompts, and its demonstrated robustness in structured decision-making. These characteristics make it particularly suitable for GUI testing, where accurate comprehension of widget semantics, layout hierarchy, and visual cues is essential for generating meaningful scenario-driven actions. The parameter temperature is fixed at 0 to promote determinism and minimize stochastic response variation across runs, ensuring consistency in logical reasoning and decision outcomes while maintaining fairness and reproducibility in testing results.
Table 1. Experiment Subject: Scenarios and Apps
Scenario Description # Task
Email Create, edit and send an email 6
Music Search, choose and plat the music 9
Note Create, edit and save the note 16
Photo Take a photo 9
Translation Choose a language, edit the text, and translate 13
Shopping Search for an item and save to the shopping cart 6
Login input the required token to login 7
Calculation Math calculation with calculator 13
Weather Query the weather of specific areas 10
Alarm-clock Create and configure the alarm-clock 10
Scenario and App Selection. To establish realistic and representative testing targets, we conduct an investigation of the Google Play Store to identify the most prevalent app categories based on download volume and the number of available apps. From these categories, we extract the most frequent and essential business functionalities as candidate testing scenarios. The test scenarios are created by three experienced software testing researchers, each with over five years of experience in mobile app testing and GUI automation. To reduce subjectivity, the researchers follow a structured scenario definition protocol: (1) each scenario is first described in natural language to represent a specific end-user goal; (2) all draft scenarios are reviewed by the research team to ensure coverage across different functional domains; and (3) overlapping or ambiguous scenarios are merged or refined through discussion. This collaborative design process ensures that the scenarios are representative of realistic user interactions while minimizing developer bias.
As summarized in Table 1, ten representative testing scenarios are finalized, covering a range of functional complexities. The first column lists the scenario names, the second provides textual descriptions, and the third shows the number of apps associated with each. Based on these scenarios, we select a total of 92 apps (with 7 used in two scenarios) and construct 99 tasks (i.e., one specific scenario on one specific app). The full list of apps is available on our project website. To enhance coverage and realism, we include both open-source and commercial apps collected from multiple major platforms, including the Google Play Store and the Huawei AppGallery. Since our LLM performs best on English-language content, only apps with English interfaces are selected to eliminate confounding effects due to multilingual GUIs. The chosen apps span a wide range of functionalities and interaction complexities, ensuring the diversity and validity of our evaluation.
Baselines. We employ several baselines to facilitate fair and meaningful comparison across research questions. For RQ1 and RQ5, we compare against three state-of-the-art scenario-based or context-aware testing frameworks: GPTDroid (Liu et al., 2024a), ScenTest (Yu et al., 2024a), and ASOT [^1] Not an official name from the authors. (Su et al., 2025). These baselines represent recent advances in scenario-guided test generation. For RQ2, which examines bug detection effectiveness, we benchmark against four representative automated GUI testing tools: Monkey (mon, 2024) (random-based), Stoat (Su et al., 2017) (model-based), PIRLTest (Yu et al., 2024b) (learning-based), and ScenTest (Yu et al., 2024a) (scenario-based), covering a spectrum of exploration strategies. This selection ensures that performance is compared across both traditional and modern paradigms in automated GUI testing. For RQ3 and RQ4, which focus on internal mechanisms of ScenGen, we use quantitative evaluations without external baselines.
4.2. RQ1: Scenario-based Test Generation Capability
In RQ1, we evaluate the overall capability of ScenGen to automatically generate GUI tests that accurately and reliably complete predefined testing scenarios. This assessment aims to verify whether ScenGen can effectively understand, plan, and execute multi-step scenario-specific testing processes, thereby demonstrating its practical usability and reliability.
To characterize the outcome of each test generation task, we classify results into four cases:
  • Case 1 ($c_{1}$): The test generation completes successfully and fully covers the target scenario.
  • Case 2 ($c_{2}$): The test generation terminates abnormally but still fully covers the target scenario.
  • Case 3 ($c_{3}$): The test generation completes successfully but fails to fully cover the target scenario.
  • Case 4 ($c_{4}$): The test generation terminates abnormally and fails to fully cover the target scenario.
Based on these classifications, we define two quantitative evaluation metrics: Coverage Rate ($C$) and Success Rate ($S$). $C$ measures the proportion of test generations that achieve complete scenario coverage, regardless of termination status, while $S$ evaluates the proportion of tests that both complete successfully and achieve full coverage. These metrics are calculated as follows:
(7) $C=\frac{|c_{1}|+|c_{2}|}{|c_{1}|+|c_{2}|+|c_{3}|+|c_{4}|}\times 100\%,\quad S=\frac{|c_{1}|}{|c_{1}|+|c_{2}|+|c_{3}|+|c_{4}|}\times 100\%$
where $\lvert c_{i}\rvert$ denotes the number of instances in case $c_{i}$.
Table 2. Result of RQ1: Scenario-based Test Generation Capability (in total of 99 tasks)
GPTDroid ScenTest ASOT ScenGen
$|c_{1}|$ 68 69 73 84
$|c_{2}|$ 6 8 11 2
$|c_{3}|$ 9 11 8 7
$|c_{4}|$ 16 11 7 6
C 74.75% 77.78% 84.85% 86.87%
S 68.69% 69.70% 73.74% 84.85%
The experimental results are summarized in Table 2. Across 99 total test generation tasks, ScenGen achieves the highest Coverage Rate (86.87%) and Success Rate (84.85%) among all compared methods. GPTDroid, ScenTest, and ASOT obtain Coverage Rates of 74.75%, 77.78%, and 84.85%, respectively, and Success Rates of 68.69%, 69.70%, and 73.74%. This comparison demonstrates that ScenGen surpasses all baselines in both comprehensive coverage and stable completion across diverse testing scenarios. Specifically, GPTDroid and ScenTest, though both scenario-oriented, rely primarily on text-based reasoning or pre-defined rules for test planning, leading to higher rates of premature or incomplete executions. ASOT achieves relatively strong results due to its adaptive scenario graph modeling, yet it lacks the semantic flexibility to handle unexpected GUI transitions or dynamically generated elements. In contrast, ScenGen benefits from its LLM-guided multi-agent design and context memory integration, allowing it to interpret GUI semantics and dynamically adapt decision-making to interface feedback.
Moreover, ScenGen exhibits notably fewer abnormal terminations (Case 2 and Case 4), indicating that the combination of multi-modal LLM reasoning and the feedback mechanism of Supervisor enables more stable iterative exploration. The smaller proportion of incomplete coverage cases (Case 3) also demonstrates that the scenario understanding of ScenGen allows it to maintain logical coherence throughout test generation, avoiding early exits that commonly affect baseline tools. From the 99 test generation tasks, ScenGen records 84, 2, 7, and 6 instances of $c_{1}$, $c_{2}$, $c_{3}$, and $c_{4}$, respectively. These results highlight that even in complex testing environments involving different app architectures and UI layouts, ScenGen can consistently generate tests that fully achieve the scenario objectives in most cases. The high coverage and success rates reflect both the reliability and generalizability of the approach across different domains.
Despite its strong overall performance, we conduct a detailed analysis of the failure cases (from $c_{2}$, $c_{3}$, and $c_{4}$) to identify improvement opportunities. The main causes of failure include premature termination, inaccurate completion detection, adaptation issues to app-specific designs, imprecise widget localization, and insufficient flexibility in handling unconventional user flows.
Premature Termination. For instance, in the “Translation” scenario, the test generation halted after text input but before triggering the final “Translate” action. Such cases are caused by transient LLM hallucinations or incomplete reasoning chains that interrupt the logical progression toward the scenario’s endpoint.
Inaccurate Completion Detection. In some “Login” scenarios, the Supervisor misjudged the successful completion of a login operation due to minimal or ambiguous visual confirmation on the GUI (e.g., lack of confirmation messages or page transitions). This highlights the dependency of the verification process on clear visual feedback, a challenge shared by both human testers and automated systems.
App-specific Design Variations. In the “Note” scenario, some apps save content implicitly upon returning from the edit page without a visible “Save” button. Such implicit operations caused ScenGen to misalign with the app’s true functional logic, reflecting the difficulty of generalizing across heterogeneous UI patterns.
Widget Localization Errors. In “Weather” apps, highly cluttered layouts led the Decider to misidentify the target search icon among visually similar icons (e.g., location or menu icons). This limitation arises from residual noise in vision-based recognition, suggesting opportunities for improved prompt engineering and fine-tuning of visual reasoning models.
Handling Unconventional Processes. In “Shopping” scenarios, when an item was out of stock, the Decider attempted to use a “find similar items” function but failed to select an alternative product and complete the purchase flow. This case demonstrates that further enhancements in contextual reasoning and adaptive planning could improve the framework’s flexibility in managing edge cases.
Overall, the results demonstrate that ScenGen outperforms existing scenario-based GUI testing frameworks across all major evaluation dimensions. Its integration of multi-modal LLM reasoning, context memory, and feedback-driven self-correction enables it to dynamically adapt to complex UI states and maintain scenario coherence throughout the testing process. Although certain failure cases remain, primarily involving ambiguous GUI feedback or app-specific design nuances, ScenGen consistently achieves higher scenario coverage, greater stability, and fewer abnormal terminations than all baseline approaches, establishing its superiority in scenario-based automated GUI test generation.
4.3. RQ2: Bug Detection Capability
To evaluate the bug detection effectiveness of ScenGen, we compare it against both traditional and scenario-based automated GUI testing approaches. For traditional methods: Monkey (mon, 2024), Stoat (Su et al., 2017), and PIRLTest (Yu et al., 2024b), we follow common experimental practice by executing each approach for two hours per task and recording all detected crashes. For scenario-based approaches: GPTDroid (Liu et al., 2024a), ASOT (Su et al., 2025), ScenTest (Yu et al., 2024a), and ScenGen, we do not impose a fixed time limit, since these methods are designed to terminate automatically upon completing the target scenario exploration.
Table 3. Result of RQ2: Bug Detection Capability
Scenario Task Monkey Stoat PIRLTest GPTDroid ScenTest ASOT ScenGen
ALL SCN ALL SCN ALL SCN
Email 6 10 3 15 7 20 3 2 2 8 9
Music 9 15 2 21 3 29 1 2 1 5 6
Note 16 25 5 35 4 50 11 8 12 15 16
Photo 9 15 4 21 3 29 2 4 5 3 6
Translation 13 21 4 29 5 41 4 10 11 13 15
Shopping 6 10 1 15 2 20 6 6 9 7 9
Login 7 12 3 17 3 23 4 2 7 9 10
Calculation 13 21 3 29 9 41 5 2 3 7 9
Weather 10 16 2 23 3 32 10 16 16 17 18
Alarm-clock 10 16 3 23 4 32 6 14 6 5 8
Sum 99 159 29 228 43 317 51 63 71 89 106
We identify detected bugs by analyzing runtime logs and filtering error messages that match the timestamps of the executed test operations, ensuring that each reported issue corresponds to an actual app failure. The results are summarized in Table 3. For traditional automated testing tools, we manually classify bugs into two categories: those occurring within the predefined testing scenarios (marked as “SCN” in Table 3) and the total number of detected bugs across the entire app (marked as “ALL”). For scenario-based approaches, all detected bugs are inherently within the target scenarios, as their exploration is scenario-constrained. All recorded bugs are manually confirmed as crash bugs. Moreover, 83.02% of the bugs detected by ScenGen have been verified or fixed by the corresponding app developers, confirming the practical relevance of the findings.
Several important observations can be drawn from the results. First, apps with higher development maturity, especially popular commercial ones, tend to exhibit fewer bugs in their core functional scenarios. Likewise, simple and heavily tested functionalities, such as “Calculation” or “Note” operations, yield fewer bug discoveries due to their inherent stability and limited behavioral complexity. In contrast, more complex business scenarios such as “Translation”, “Shopping”, and “Login” produce significantly higher bug counts, where multi-step interactions and asynchronous GUI transitions are more prone to runtime errors. Manual revalidation confirms that all apps in scenarios with no detected bugs indeed operated correctly within their target functionality.
In comparing bug detection across different approaches, several trends are evident. Traditional automated testing methods detect more bugs in total (“ALL”), as they explore the entire app without scenario constraints. However, within specific target scenarios, their coverage and precision are notably lower than those of scenario-based methods. In contrast, ScenGen demonstrates strong scenario fidelity and high bug detection rates across nearly all functional categories. Importantly, upon manual verification, we confirm that all bugs detected by any baseline approach are also detected by ScenGen. No baseline tool identifies a unique bug that ScenGen fails to discover. This result underscores the comprehensiveness of exploration of ScenGen and its ability to cover both shallow and deep interaction paths relevant to the testing scenarios.
The superior bug detection capability of ScenGen can be attributed to several factors. First, guided by the LLM’s semantic understanding, ScenGen can open hidden or collapsed menus to uncover target widgets that are otherwise inaccessible to random or model-based tools. This behavior allows ScenGen to reach interaction paths that directly contribute to fulfilling the scenario goals. In contrast, traditional testing tools tend to abandon such paths once obstacles are encountered, diverting to unrelated functionalities. Second, the test operation sequences generated by ScenGen are typically longer and semantically coherent, as the LLM-guided reasoning ensures that each step advances the scenario progression. This deeper exploration enables the detection of complex bugs that only manifest after multiple dependent interactions, which are often missed by shorter or random sequences generated by traditional approaches. Finally, the multi-agent design ensures that erroneous operations are detected and corrected in real time through the Supervisor, maintaining test traceability and consistent alignment with the target scenario.
Overall, ScenGen exhibits strong bug detection capabilities across diverse app types and testing scenarios. Its LLM-guided and scenario-focused exploration enables it to uncover distinct and meaningful bugs that traditional approaches either overlook or cannot reproduce consistently. The fact that every bug detected by baseline approaches is also detected by ScenGen highlights its superior coverage and robustness. Consequently, ScenGen not only complements existing automated testing strategies but also establishes itself as an effective framework for precise and comprehensive scenario-based GUI testing.
4.4. RQ3: Widget Localization Effectiveness
Widget localization plays a crucial role in enabling accurate and reliable GUI interactions during automated testing. Unlike pure vision-based or text-based methods, ScenGen integrates the pixel-level precision of traditional computer vision (CV) algorithms with the semantic reasoning capabilities of multi-modal LLMs to handle visually complex and context-rich user interfaces. RQ3 aims to evaluate the effectiveness of this hybrid strategy in widget matching and localization, examining how the integration of CV techniques and LLM-driven understanding improves both accuracy and robustness in real-world testing environments.
We measure localization accuracy as the key evaluation metric, capturing both the initial localization performance and the improvement after self-correction. Specifically, the initial localization accuracy $Acc_{1}$ and final localization accuracy $Acc_{f}$ are computed as follows:
(8) $Acc_{1}=\frac{n_{1}}{N}\times 100\%,\quad Acc_{f}=\frac{n_{f}}{N}\times 100\%,\quad\uparrow=\frac{Acc_{f}-Acc_{1}}{Acc_{1}}\times 100\%$
where $n_{1}$ represents the number of successful widget localizations without correction, $n_{f}$ denotes the number of final successful localizations after the correction process, and $N$ is the total number of localization attempts.
For ScenGen, widget localization and correction are considered a continuous, feedback-driven process inspired by human visual interaction. When an initial localization fails due to visual ambiguity, overlapping elements, or partial visibility, the Decider triggers the self-correction mechanism, which re-analyzes the app state using both visual feedback and semantic reasoning. A localization attempt is deemed successful if the subsequent operation is executed correctly and produces the expected GUI state transition. The correctness of each localization result is independently reviewed by three evaluators, each with over five years of mobile app testing experience. Evaluators assess localization success solely based on the app GUI and operational outcome, without knowledge of system internals or other evaluations. A localization is confirmed as correct only when all three judges unanimously agree, ensuring consistency and minimizing subjectivity in assessment.
Table 4. Result of RQ3: Widget Localization Effectiveness
Scenario N $n_{1}$ $n_{f}$ $Acc_{1}$ $Acc_{f}$ $\uparrow$
Email 33 28 32 84.85% 96.97% 14.29%
Music 22 12 22 54.55% 100.00% 83.33%
Note 53 44 52 83.02% 98.11% 18.18%
Photo 9 6 9 66.67% 100.00% 50.00%
Translation 35 34 34 97.14% 97.14% 0.00%
Shopping 31 26 30 83.87% 96.77% 15.38%
Login 35 32 33 91.43% 94.29% 3.13%
Calculation 44 34 44 77.27% 100.00% 29.41%
Weather 27 21 26 77.78% 96.30% 23.81%
Alarm-clock 23 21 23 91.30% 100.00% 9.52%
Sum/Avg 312 258 305 80.79% 97.96% 24.71%
The experimental results are summarized in Table 4. Across ten functional scenarios, the initial localization accuracy ($Acc_{1}$) ranges from 54.55% to 97.14%, while the final localization accuracy ($Acc_{f}$) after self-correction improves substantially, reaching between 94.29% and 100%. In four scenarios, the final accuracy achieves a perfect 100%. On average, the initial accuracy is 80.79%, and after correction, the average rises to 97.76%, with all scenarios over 94% in the final stage.
This improvement clearly demonstrates the effectiveness of the self-correction mechanism and the synergy between traditional CV-based detection and LLM-based semantic interpretation. The correction process allows the system to identify and recover from initial misalignments caused by dynamic layouts or ambiguous widget boundaries. In nine out of ten scenarios, accuracy increases following correction, with gains ranging from 3.13% to 83.33%. These results confirm that ScenGen not only maintains high localization precision in stable UI conditions but also exhibits strong fault tolerance under dynamically changing or visually cluttered interfaces.
Furthermore, qualitative analysis reveals that most initial localization errors occur in cases where widgets are visually occluded (e.g., drop-down or scrollable menus) or contain highly similar visual components (e.g., repeated icons). The self-correction mechanism successfully resolves these issues by leveraging LLM-based semantic reasoning to re-evaluate contextual cues and by refining coordinate-level predictions based on CV feedback. This cooperative process effectively mimics human testers’ behavior when re-targeting misclicked or ambiguous GUI elements. Overall, RQ3 demonstrates that combining computer vision algorithms with multi-modal LLM reasoning significantly enhances widget localization reliability in GUI testing. The self-correction mechanism further strengthens this process by dynamically adapting to interface changes and mitigating recognition errors. These results validate that ScenGen achieves precise, context-aware localization across diverse GUI layouts, ensuring a stable foundation for accurate and scenario-driven automated test execution.
4.5. RQ4: Logical Decision-Making Effectiveness
RQ4 evaluates whether the Decider can accurately interpret the requirements of a testing scenario and generate appropriate next-step operations based on the current app GUI state. The goal is to assess both the logical soundness of initial decisions and the effectiveness of the self-correction mechanism when feedback from the environment necessitates revision. We use decision accuracy as the primary evaluation metric. To measure the impact of the self-correction mechanism, each decision and its potential subsequent correction are treated as a single decision unit. We then calculate the initial accuracy ($Acc_{1}$) and final accuracy ($Acc_{f}$) according to:
(9) $Acc_{1}=\frac{n_{1}}{N}\times 100\%,\quad Acc_{f}=\frac{n_{f}}{N}\times 100\%,\quad\uparrow=\frac{Acc_{f}-Acc_{1}}{Acc_{1}}\times 100\%$
where $n_{1}$ denotes the number of initially correct decisions, $n_{f}$ the number of finally correct decisions after self-correction, and $N$ the total number of decisions in the given testing scenario.
When recording decisions, we treat both normal operation selections and test completion decisions as decision units. A decision is considered correct if the action it recommends advances the test scenario toward its expected goal or terminates it appropriately once the goal is achieved. To ensure objectivity and consistency, three independent evaluators serve as judges, all with more than five years of professional experience in mobile app testing. Each evaluator receives only the scenario description, GUI state, and the LLM’s decision, without access to the internal reasoning process or other evaluators’ results. “Common expectations” are defined as the expected user actions that a typical human tester would perform to complete the scenario goal, based on the scenario definition established before the experiments. For instance, in a “Login” scenario, entering valid credentials and submitting the form is considered the expected path; in a “Shopping” scenario, inputting a query and confirming submission qualify. A decision is counted as correct only when all three evaluators unanimously agree that the generated operation aligns with the defined scenario intent. This majority-independent assessment mitigates bias, even though the authors contributed to the original scenario design.
The number of decisions recorded for each test scenario represents the total number of action decisions made across all apps associated with that scenario. For example, the 39 decisions listed under the “Email” scenario in Table 5 correspond to the aggregate decisions generated from all six tasks related to that functional scenario (as shown in Table 1).
Table 5. Result of RQ4: Logical Decision-Making Effectiveness
Scenario N $n_{1}$ $n_{f}$ $Acc_{1}$ $Acc_{f}$ $\uparrow$
Email 39 38 39 97.44% 100.00% 2.63%
Music 31 31 31 100.00% 100.00% 0.00%
Note 71 64 66 90.14% 92.96% 3.13%
Photo 18 18 18 100.00% 100.00% 0.00%
Translation 51 48 48 94.12% 94.12% 0.00%
Shopping 38 36 36 94.74% 94.74% 0.00%
Login 42 40 41 95.24% 97.62% 2.50%
Calculation 57 57 57 100.00% 100.00% 0.00%
Weather 28 27 27 96.43% 96.43% 0.00%
Alarm-clock 33 33 33 100.00% 100.00% 0.00%
Sum/Avg 408 392 396 96.08% 97.06% 1.02%
Across the ten evaluated scenarios, the initial decision accuracy ranges from 90.14% to 100.00%, while the final accuracy ranges from 92.96% to 100%. On average, the initial accuracy ($Acc_{1}$) is 96.82%, and the final accuracy ($Acc_{f}$) increases to 97.59%, demonstrating that the self-correction mechanism yields a measurable improvement in decision reliability. In four scenarios, the Decider achieves perfect accuracy without correction, whereas in three other scenarios, accuracy improves noticeably after correction. This result confirms that the self-correction mechanism effectively compensates for transient decision errors caused by dynamic GUI updates or ambiguous widget states.
A detailed examination of the 12 incorrect decisions reveals distinct error types. Six errors stem from premature test termination before completing the full scenario (e.g., stopping after inputting credentials but before confirming submission). Two cases result from redundant continuation after the scenario goal has already been satisfied. Among the remaining four cases, two involve misinterpretation of GUI widget semantics, such as confusing similarly labeled buttons (“Next” vs. “Submit”), while one error arises from missing prerequisite operations (e.g., failing to open a required menu before proceeding). Another error occurs when the test sequence enters a rare alternative path that deviates from the normal scenario flow. Notably, four operational errors are successfully corrected by the self-correction mechanism, which re-evaluates the app state and adjusts the decision accordingly. This demonstrates adaptive capability of ScenGen to recover from suboptimal decisions through iterative feedback.
In summary, the Decider exhibits high decision accuracy and strong adaptability across diverse app scenarios. The inclusion of a self-correction mechanism ensures robustness against dynamic GUI changes and reasoning inconsistencies. These findings indicate that the LLM-guided decision-making process in ScenGen effectively captures the logical intent of scenario-based testing and reliably translates it into executable GUI operations.
4.6. RQ5: Scenario-based Test Generation Efficiency
The efficiency of automated test generation is crucial for ensuring the practicality of GUI testing frameworks. However, for LLM-based systems, runtime can be highly influenced by external factors such as API response latency and network stability. Consequently, we adopt a two-fold evaluation strategy focusing on both token consumption and execution time overhead. The former provides a model-independent measure of computational demand, while the latter quantifies overall testing efficiency in real-world deployments.
4.6.1. Token Consumption Analysis
To accurately assess the operational efficiency of ScenGen, we analyze its token usage across different testing scenarios. Specifically, we record the number of tokens consumed by each API call, categorized into six stages: logical decision-making ($T_{1}$), widget localization ($T_{2}$), loading state verification ($T_{3}$), app GUI state verification ($T_{4}$), test completion verification ($T_{5}$), and self-correction analysis ($T_{6}$). The total token consumption for each scenario ($T_{total}$) is also computed as the sum across all stages.
Table 6. Result of RQ5: Token Overhead of Scenario-based Test Generation
Scenario $T_{1}$ $T_{2}$ $T_{3}$ $T_{4}$ $T_{5}$ $T_{6}$ $T_{total}$
Email 5586 8295 5281 13970 11823 833 45788
Music 3748 10197 4982 14729 8841 1963 44460
Note 6461 13639 6602 21667 12178 2573 63120
Photo 1486 3057 1791 6533 3571 746 17184
Translation 4645 6452 5640 11758 10296 359 39150
Shopping 8782 14461 12845 26102 19164 2757 84111
Login 8434 13263 7830 21329 18205 1253 70314
Calculation 5167 9184 6116 15603 12117 1940 50127
Weather 4522 11104 5474 12961 9419 1696 45176
Alarm-clock 3490 3794 3229 9120 8270 456 28359
Average 5232 9345 5979 15377 11388 1458 48779
Table 6 summarizes the token consumption in different scenarios, revealing a wide range from 17,184 to 84,111 tokens. Complex scenarios such as “Shopping” and “Login” consume more tokens due to their intricate, multi-step interactions and the frequent involvement of the self-correction mechanism. In contrast, simpler scenarios like “Photo” and “Alarm-clock” require fewer tokens because they involve fewer decision branches and minimal contextual reasoning.
A closer examination of the six stages indicates that the app GUI state verification ($T_{4}$) and test completion verification ($T_{5}$) stages generally exhibit the highest token consumption. This is expected, as these stages require multi-page visual comparisons and integrate both semantic reasoning and visual understanding to determine test progression or termination. Conversely, logical decision-making ($T_{1}$) and self-correction analysis ($T_{6}$) consume fewer tokens because they typically process less visual input. Widget localization ($T_{2}$) and loading state verification ($T_{3}$) show moderate consumption, reflecting their balanced use of visual and textual information.
Overall, the token usage results highlight that the computational cost of ScenGen aligns with scenario complexity and interaction richness. The modular token distribution across different stages also confirms the rational design of the agent collaboration workflow, where resources are concentrated on the most information-intensive reasoning tasks.
4.6.2. Time Overhead Analysis
In addition to token efficiency, we evaluate the time overhead of ScenGen in comparison with other state-of-the-art scenario-based testing approaches, including GPTDroid (Liu et al., 2024a), ScenTest (Yu et al., 2024a), and ASOT (Su et al., 2025). We measure the average test generation time (in minutes) per scenario, calculated from multiple runs to ensure stability.
Table 7. Result of RQ5: Time Overhead of Scenario-based Test Generation (minute)
Scenario GPTDroid ScenTest ASOT ScenGen
Email 6.77 1.27 8.46 4.22
Music 7.13 1.24 7.63 3.37
Note 7.21 1.94 12.37 4.65
Photo 8.97 2.95 13.26 5.86
Translation 7.37 1.35 8.16 4.65
Shopping 6.58 2.16 11.69 4.81
Login 8.71 1.82 12.70 5.55
Calculation 7.36 2.10 10.68 3.15
Weather 6.78 1.30 11.54 5.50
Alarm-clock 7.93 1.05 7.35 5.39
Average 7.48 1.72 10.38 4.71
As shown in Table 7, the average time overhead of ScenGen (4.71 minutes) is higher than that of ScenTest (1.72 minutes) but significantly lower than that of ASOT (10.38 minutes) and GPTDroid (7.48 minutes). This demonstrates that ScenGen achieves a balanced trade-off between test generation efficiency and scenario reasoning capability. While its runtime exceeds ScenTest due to additional LLM interactions and vision-language inference, the increased computational effort translates directly into superior coverage and success rates, as demonstrated in RQ1.
Examining per-scenario results, we observe that ScenGen maintains stable time consumption across different application domains, with smaller variations than the baselines. For example, in complex scenarios such as “Shopping” and “Login”, ASOT and GPTDroid exhibit substantial time inflation (above 11 minutes on average), whereas ScenGen completes comparable tasks within 4–5 minutes. This stability stems from the adaptive decision control and dynamic context management of ScenGen, which prevent redundant explorations and minimize backtracking operations.
RQ5 confirms that ScenGen achieves efficient and resource-aware scenario-based test generation. Its token consumption scales proportionally with scenario complexity, reflecting effective resource allocation across different reasoning stages. In terms of time overhead, ScenGen demonstrates moderate runtime with stable performance across diverse app scenarios, significantly faster than other advanced LLM-based frameworks while maintaining superior reliability and adaptability. These findings affirm that ScenGen provides a computationally efficient yet semantically powerful solution for real-world scenario-based GUI testing.
5. Threats to Validity
To ensure the rigor and credibility of our experimental evaluation, we analyze potential threats to validity that may influence the interpretation or generalization of our results. The main threats involve the selection of scenarios and apps, the stochastic behavior of the LLMs, and the possibility of LLM hallucinations during testing. For each identified threat, we discuss its potential impact and the corresponding mitigation strategies adopted in ScenGen.
Selection of Scenarios and Apps. A potential threat to validity arises from the limited number and the potentially arbitrary selection of app scenarios used in the experiments. To mitigate this threat, we adopt a systematic scenario selection process that balances both diversity and representativeness. Specifically, we include testing scenarios that reflect common functionalities across multiple domains of daily mobile usage (e.g., communication, productivity, entertainment, and utilities), while also ensuring variation in interaction complexity and interface design. The selected scenarios were reviewed and refined by multiple developers and testers to reduce personal bias. Furthermore, the experimental dataset includes a mix of popular open-source and commercial apps collected from multiple sources, ensuring that the evaluation covers heterogeneous app architectures and design styles. These measures collectively enhance the representativeness, breadth, and practical relevance of the experimental results.
Instability of LLMs. The probabilistic nature of the LLMs may cause instability when interpreting complex contextual inputs, leading to variations in outputs even when the temperature parameter is fixed at 0. To minimize this risk, we conduct repeated runs of each test scenario and report aggregated results obtained over 10 repetitions, rounded to the nearest integer. This repetition-based strategy filters out random variations and highlights stable behavioral patterns, thereby improving the reliability of the reported results and the consistency of agent behavior during test generation.
LLM Hallucinations. Another potential threat lies in LLM hallucinations, where the model may generate logically plausible but practically invalid actions, such as interacting with non-existent widgets or misidentifying visual elements. To address this issue, ScenGen introduces the supervisory role of the Supervisor. After each action execution, the Supervisor verifies whether the resulting app state aligns with the intended scenario. When deviations are detected, it initiates a rollback, reanalyzes the decision, and executes a corrected action. This closed-loop supervision ensures that each generated action remains consistent with both the target scenario and the real GUI state, thus mitigating the adverse impact of hallucinations on test validity.
6. Related Work
6.1. Vision-based GUI Testing
Recent advancements in vision-based GUI testing have significantly enhanced both automated and manual testing processes by enabling a more intuitive understanding of the app GUI through visual information (YazdaniBanafsheDaragh and Malek, 2021). This research direction leverages computer vision (CV), deep learning, and multimodal analysis to perceive and reason about GUI layouts, thereby addressing the limitations of traditional structure- or code-based approaches.
Early efforts primarily focus on visual defect detection and widget recognition. Liu et al. (Liu et al., 2020) propose OwlEye, a tool that detects display inconsistencies in GUI screenshots. White et al. (White et al., 2019) apply machine learning techniques to identify GUI widgets, thereby improving the efficiency of random testing strategies. Xie et al. (Xie et al., 2020) introduce UIED, which integrates traditional computer vision with deep learning to enhance widget detection accuracy, while Zhang et al. (Zhang et al., 2021) develop a more efficient model for GUI element detection. Mansur et al. (Hasan Mansur et al., 2023) present AidUI, which combines CV and NLP techniques to detect dark patterns, improving the ethical and usability evaluation of user interfaces. Besides, Baral et al. (Baral et al., 2024) automate the GUI-based test oracle construction for mobile apps, further illustrating the potential of vision-driven approaches in improving testing reliability.
As vision-based understanding matures, subsequent studies integrate visual perception with automated test generation and cross-platform analysis. Yu et al. (Yu et al., 2024b) combine deep image understanding with reinforcement learning to develop a cross-platform testing framework, demonstrating that semantic understanding of GUI screenshots can guide more intelligent exploration. Zhang et al. (Zhang et al., 2023) propose ReSPlay, a record-and-replay framework enhanced with visual features to improve replay robustness. Ji et al. (Ji et al., 2023) investigate vision-based widget mapping to facilitate cross-platform test migration, while Yu et al. (Yu et al., 2024a) construct event knowledge graphs from GUI images to guide scenario-based testing, highlighting the potential of visual semantics for bridging low-level interface elements with high-level functional intent.
Beyond automation, vision-based methods have also improved the efficiency of manual testing and maintenance tasks. Fazzini and Orso (Fazzini and Orso, 2017) develop DiffDroid to detect cross-platform GUI inconsistencies. Wang et al. (Wang et al., 2019) automate the identification and elimination of duplicate test reports, while Xu et al. (Xu et al., 2021) introduce GUIDER to repair Android GUI test scripts. Similarly, Yoon et al. (Yoon et al., 2022) address View Identification Failures in smoke tests, aiding smartphone vendors in maintaining test scripts across evolving app versions. Yu et al. (Yu et al., 2022) propose automatic bug inference and description generation from GUI test reports, and Yan et al. (Yan et al., 2024) improve duplicate detection in video-based bug reports using vision transformers, demonstrating the growing influence of deep visual representations.
Overall, the integration of visual data from GUI screenshots has substantially improved the accuracy, adaptability, and coverage of GUI testing. By enabling visual-semantic understanding of app interfaces, these techniques bridge the gap between low-level widget detection and higher-level functional reasoning. However, most existing approaches remain limited to perception and pattern recognition, lacking the ability to reason about testing scenarios and business logic. This gap motivates our work on ScenGen, which extends vision-based GUI understanding with LLM-guided semantic reasoning, enabling context-aware and scenario-driven automated testing.
6.2. LLMs for Software Testing
The rapid advancement of LLMs (Devlin et al., 2019; Ouyang et al., 2024) has significantly influenced diverse research domains, including natural language understanding, reasoning, and multimodal learning (Kojima et al., 2022; Wei et al., 2024; Brown et al., 2020). Their powerful generalization and reasoning capabilities have made them particularly promising for complex software engineering tasks, where semantic comprehension and context awareness are essential. In the field of software testing, the application of LLMs to automate and enhance various testing activities has recently emerged as a compelling direction of research.
For GUI testing, in particular, LLMs offer the potential to reason about user intentions and testing scenarios, thereby bridging the gap between low-level widget interactions and high-level business logic (Yang et al., 2023). This perspective opens new possibilities for scenario-driven test generation, where LLMs can infer the semantics of GUI elements, understand interaction contexts, and produce meaningful test sequences aligned with functional goals, capabilities that traditional rule- or model-based methods struggle to achieve.
Recent studies have explored different dimensions of applying LLMs to software testing (Jalil et al., 2023; Fu et al., 2022). In test case generation, Dakhel et al. (Dakhel et al., 2024) enhance LLM-generated tests using mutation testing to improve diversity and fault detection. Chen et al. (Chen et al., 2024) propose ChatUniTest, an LLM-based framework for automated unit test generation, while Schafer et al. (Schäfer et al., 2024) conduct a large-scale empirical study evaluating the effectiveness of LLMs for automated unit test creation. Similarly, Yu et al. (Yu et al., 2023) leverage LLMs to generate mobile app test scripts, demonstrating their capacity to understand app contexts and produce executable GUI-level actions.
In GUI-specific testing, LLMs have shown promising abilities in understanding visual and contextual cues. Liu et al. (Liu et al., 2023) use LLMs to generate context-appropriate textual inputs, while Liu et al. (Liu et al., 2024a) formulate mobile GUI testing as a question–answering task, significantly improving coverage and generalization. Another study by Liu et al. (Liu et al., 2024b) employs LLMs to generate rare or extreme input values, effectively identifying app crashes. These works illustrate how natural language reasoning enables more meaningful and scenario-aware test generation for GUI-based systems.
Beyond test generation, LLMs have been applied to bug detection and program repair. Huang et al. (Huang et al., 2024) utilize pre-trained LLMs for crash reproduction based on stack traces, and Feng et al. (Feng and Chen, 2024) propose a lightweight prompting technique to reproduce bugs from textual reports. In the broader area of program repair, Jin et al. (Jin et al., 2023) introduce a fine-tuned LLM framework for defect localization and patch generation, while Ribeiro et al. (Ribeiro et al., 2022) employ CodeGPT for automated program repair. Pearce et al. (Pearce et al., 2023) further analyze the zero-shot vulnerability repair capability of LLMs, showing that modern models can capture implicit program semantics even without explicit supervision.
Collectively, these studies demonstrate the transformative potential of LLMs in software testing, improving efficiency, accuracy, and semantic understanding across various testing tasks. With the continued evolution of multi-modal LLMs, capable of processing text, images, and structured data, their integration into GUI testing offers unprecedented opportunities. By combining visual perception with scenario reasoning, LLMs can move beyond static event exploration toward context-aware, human-like test generation. This insight underpins the motivation for our proposed approach, ScenGen, which employs LLM-guided, scenario-based reasoning to enhance the completeness and intelligence of automated GUI testing.
7. Conclusion
The rapid evolution of mobile apps has imposed increasingly stringent requirements on software quality assurance. Although numerous automated GUI testing approaches have been developed, they still face fundamental challenges, including limited flexibility in adapting to diverse app interfaces and an insufficient understanding of app functionality scenarios. These limitations hinder existing methods from fully capturing realistic user interactions and comprehensively validating app behavior. To address these challenges, this paper introduced ScenGen, a novel LLM-guided scenario-based GUI testing framework. ScenGen leverages multi-modal LLMs to comprehend app semantics and testing scenario intent, while employing a multi-agent collaboration architecture to simulate the human testing process. The framework is structured into five coordinated stages, managed respectively by the Observer, Decider, Executor, Supervisor, and Recorder agents, supported by a shared context memory. This modular design ensures that perception, decision-making, execution, verification, and recording are clearly delineated, enhancing both the traceability and adaptability of the testing process. Experimental results demonstrate that ScenGen effectively generates scenario-consistent GUI tests, capturing the semantic relationships between GUI elements and user goals. These findings highlight the potential of integrating LLM-driven reasoning and multi-agent collaboration to advance automated GUI testing toward human-like, scenario-aware testing intelligence.
In summary, ScenGen provides a new paradigm for automated GUI testing by uniting vision-based understanding, LLM-guided reasoning, and scenario-oriented test generation. Future research will explore expanding the framework to broader application domains, incorporating continual learning for cross-app generalization, and further enhancing the robustness of LLM-guided testing under dynamic GUI environments.
Acknowledgements.
The authors would like to thank the editors and anonymous reviewers for their time and comments. This work is partially supported by the National Key Research and Development Program of China (2024YFF0908005).